Jan got a test of 92%. She had 35 items correct. How many test items were there?
Answer by jim_thompson5910(35256)

x = total number of items

35/x = 92/100

35*100 = 92*x

3500 = 92x

92x = 3500

x = 3500/92

x = 38.0434782608696

x = 38 ... round to the nearest whole number

So there are 38 test questions total.